The ratio of C p /C v for inert gases is (A) 1.33 (B) 1.66 (C) 2.13 (D) 1.99



The ratio of Cp/Cv for inert gases is    


        (A)    1.33


        (B)    1.66


        (C)    2.13


(D)    1.99


Grade:upto college level

1 Answers

Simran Bhatia
10 years ago

Inert gas exist as mono atomic molecules. For these,

    C_p=5/2 R

    C_v=3/2 R

    γ=C_p/C_v =(5/2 R )/(3/2 R)=1.66